OpenStudy (anonymous):

The sum of two numbers is 17. Their difference is 45. Find the numbers. Separate your answers using a coma.

OpenStudy (anonymous):

31 and -14

OpenStudy (anonymous):

x+y=17

OpenStudy (anonymous):

the difference is 45 so x-45=17

OpenStudy (anonymous):

Let x be the the first number Let y be the second number then, x + y = 17 x - y = 45 solving the systems 2x = 62 x = 31 to determine y, insert 31 to any of the two equations y = - 14 ANS. 31, -14

OpenStudy (anonymous):

O.K.